f(x)=3x^2+4x-7 finded f(x+1)-f(x)
Answer by samuenkadi(3)

f(x)=3x^2+4x-7 finded f(x+1)-f(x)
f(x+1)=3(X+1)^2+4x-3
3x^2+10x-3

f(x+1)-f(x)=6x+4